OpenStudy (anonymous):

please help, how many 5 digit numbers can be formed from the digits 1,2,3,4,5,6,7

OpenStudy (ranga):

Can the digits repeat?

OpenStudy (anonymous):

no

OpenStudy (ranga):

The first digit can be any of the 7 numbers 1,2,3,4,5,6,7. Having used up one digit and because we are not allowed to repeat the digit, the second can be any of the six digits left. The third can be any of the five digits left, etc. So the number of possibilities are: 7 x 6 x 5 x 4 x 3 = 2520

OpenStudy (anonymous):

This is a basic statistic math class dealing with counting, combinations, and for whatever reason I am stuck on this question

OpenStudy (anonymous):

to be honest that was the first answer I had, but was unsure if it was correct

OpenStudy (ranga):

You are correct. If the digits can repeat itself the answer will be 7x7x7x7x7

OpenStudy (anonymous):

actually I think I x the 2 and 1 as well

OpenStudy (anonymous):

not the 1 but the 2, may I asked why did you take away the 2, because i did not

OpenStudy (ranga):

They are asking for a FIVE digit number: First digit: 7 possibilities Second digit: 6 possibilities Third digit: 5 possibilities Fourth digit: 4 possibilities Fifth digit: 3 possibilities Total possibilities for a FIVE digit number = 7x6x5x4x3

OpenStudy (anonymous):

ok, I understand Thank you

OpenStudy (anonymous):

Can you assist me with two more questions please? I thank you for your help, I really appreciate it

OpenStudy (ranga):

Let me try one more. My statistics is a bit rusty

OpenStudy (anonymous):

I don't know anything about poker, never been to vegas, I don't understand the card game. Explain how you would find the total number of Five card poker hands? Indicate what this number would be

OpenStudy (ranga):

I don't know the game either! Just look up online for five card poker and see what they say.

OpenStudy (anonymous):

ok, I will try that what about this question, how many groups of three can be formed using Tom, John, Mary, Sally and Sue

OpenStudy (ranga):

Do you know the formula for permutations and combinations?

OpenStudy (anonymous):

yes

OpenStudy (ranga):

\[\Large ^{n}C _{r} = \frac{ n! }{ r!(n-r)! }\]

OpenStudy (anonymous):

I k now that the order is not important in forming groups and is also not important in dealing cards

OpenStudy (anonymous):

ok so i would use Combinatorial number nCr=N! /R! (N-R)!

OpenStudy (anonymous):

so its n objects taking R at a time

OpenStudy (ranga):

Yes. So in the 5 card poker hand they are asking how many ways 5 cards can be drawn from a deck of 52 cards. n = 52, r = 5 Plug it into the formula

OpenStudy (anonymous):

I understand that part am not getting how many groups of three , my answer was 5

OpenStudy (anonymous):

ok so we are going back to the poker

OpenStudy (anonymous):

for the poker I had 52C5

OpenStudy (ranga):

In the next question you have 5 people and they are asking how many ways you can select 3 people out of the 5. n = 5, r = 3

OpenStudy (anonymous):

so would that be 5*4*3

OpenStudy (ranga):

Yes poker is 52C5 = 52!/(5!47!) = 52x51x50x49x48/(5x4x3x2x1) = ?

OpenStudy (ranga):

5C3 = 5!/(3!2!) = 5x4/2 = 10

OpenStudy (anonymous):

ok the 5C3 number of five people, question, so 5!/ (3!2!)5-3=2 where did the 4 come from

OpenStudy (anonymous):

we are going back to the question of 5 people and how many ways you can select 3 people out of five

OpenStudy (anonymous):

5C3=5!/(3!2!)

OpenStudy (ranga):

\[\Large 5C3 = \frac{ 5! }{ 3!2! } = \frac{ 5\times4\times3\times2\times1 }{ (3\times2\times1)(2\times1)}\]

OpenStudy (ranga):

See how the 3x2x1 in the numerator and the denominator cancel out?

OpenStudy (anonymous):

yes

OpenStudy (ranga):

So instead of expanding 5! all the way I knew I can stop it at 5! = 5x4x3! because I see a 3! in the denominator that will cancel it out.

OpenStudy (anonymous):

I understand, Thank You!

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

How often are you on this site for assistance?

OpenStudy (ranga):

As time permits. But don't be hesitant. Just post the questions in the general area and there are a number of people who would jump in to help you out.

OpenStudy (anonymous):

I am taking classes online and I am active duty military, so it is hard to find a math tutor at night.

OpenStudy (anonymous):

I have not taken a math class in over 10 years I really appreciate your help, have a good night

OpenStudy (ranga):

You too.